Vertex (0,1)
Focus (0,4)
what is the equation?

Answer by lwsshak3(11628)

Vertex (0,1)
Focus (0,4)
what is the equation?
***
Given data shows this is an equation of a parabola that opens up with vertex at(0,1) and focus (0,4)
Its basic equation:
(x-h)^2=4p(y-k), (h,k)=coordinates of the vertex
axis of symmetry: x=0 or y-axis
p=3 (distance from vertex to focus on the axis of symmetry)
4p=12
equation: x^2=12(y-1)